OpenStudy (anonymous):

MEDAL AND BEST RESPONSE! HELP!! Rectangle A has an area of 4 - x^2. Rectangle B has an area of x^2 + 2x - 8. In simplest form, what is the ratio of the area of Rectangle A to the area of Rectangle B? Show your work.

OpenStudy (anonymous):

\[4-x^2=-(x-2) (x+2), x^2+2 x-8=(x-2) (x+4) \]